# Relativistic corrections to energy spectrum of hydrogen due to the full one-photon-exchange interaction

### v2

@@ Abstract
In this work, we estimate the relativistic corrections to the energy spectrum of hydrogen resulting from the full one-photon-exchange interaction using a highly precise numerical method. In the frame of the effective Schrodinger-like equation, which is derived exactly from the Bethe-Salpeter equation in quantum electrodynamics, we express the effective potential corresponding to the full one-photon-exchange interaction in terms of eight scalar functions. Unlike the usual calculations performed in the references, where the effective potential is expanded in terms of momenta order by order, we retain the exact momentum dependence in the effective potential to estimate its corrections to the energy spectrum of hydrogen using a highly precise numerical method. We also use the same numerical method to double-check the results in the non-relativistic case to ensure accuracy. We discuss the comparison of the numerical results with those obtained using the usual bound-state perturbative theory. Our calculations suggest that it is possible to accurately account for all the relativistic contributions using this method. It would be interesting to extent these calculations to positronium, muonic hydrogen, and cases involving nuclear structure and radiative corrections.
@@ I Introduction
Throughout history, various other methods have also been employed to investigate corrections beyond the Coulomb potential. These include the use of the Dirac equation with an effective potential Grotch-RMP-1969 , the external field approximation Eides-Book-2007 , the quasi-potential approach Faustov-1985-TMP ; Faustov-1999-zETF ; Martynenko-2005-JETP , and the Foldy-Wouthuysen transformation method Foldy-PR-1950 ; Pachucki-2005-PRA ; Zhou-JPB-2023 , among others. However, in this work, we focus our discussion within the framework of the effective Schrodinger-like equation, which is derived exactly from the Bethe-Salpeter equation in QED Lepage-1978-PRA-Schrodinger-like-equation .
In the bound-state perturbative theory, the interaction kernel is expanded order by order in terms of the coupling constant $\alpha_{e}$, velocities $\vec{p}_{i}/m_{e,p}$, and $m_{e}/m_{p}$, where $\vec{p}_{i}$ represents the three momenta of the system, and $m_{e}$ and $m_{p}$ denote the masses of the electron and proton, respectively. The expansion of the interaction kernel in terms of momenta implies that only contributions from low momenta are accurately considered, while contributions from high momenta need to be treated separately.
On the other hand, the wave functions in the Schrodinger equation with the Coulomb potential are valid for the entire momentum range. The integration of the wave functions and interaction kernels encompasses contributions from both low and high momenta. This can introduce additional ultraviolet (UV) divergences, which differ somewhat from the usual UV divergences in the scattering amplitudes arising from loop integrals in radiative corrections. To handle these UV divergences, regularization methods are employed Pachucki-PRA-1997 ; Czarnecki-PRA-1999 ; Adkins-PRA-2020 . These UV divergences eventually cancel each other out, and the final results are independent of the regularization procedure.
In this study, we use the one-photon-exchange (OPE) interaction as an example to perform a numerical calculation of the energy spectrum. Instead of expanding the OPE interaction in terms of momenta and $m_{e}/m_{p}$, we retain it in its full form. This approach allows us to include all relativistic corrections at the order of $\alpha_{e}$, and it avoids any additional UV divergences since the full OPE interaction kernel remains valid across the entire momentum range. This calculation can be regarded as a full result incorporating the OPE interaction, and it provides an interesting basis for comparison with the results obtained using the bound-state perturbative theory.
@@ II Basic Formula
II.1 The effective Schrodinger-like equation
In QED, the bound states of the $ep$ system in the center frame can be accurately described by the following effective Schrodinger-like equation Lepage-1978-PRA-Schrodinger-like-equation :
@@ II.1 The effective Schrodinger-like equation
where $P=(P_{0},\vec{0})\equiv p_{1}+p_{2}=p_{3}+p_{4}$, $p_{1,2,3,4}$ are the momenta of final and intermediate fermions as shown in Fig.1, $E_{i}^{(e)}\equiv\sqrt{\vec{p}_{i}^{2}+m_{e}^{2}}$, $E_{i}^{(p)}\equiv\sqrt{\vec{p}_{i}^{2}+m_{p}^{2}}$. The interaction kernel $\widetilde{K}_{\bar{\lambda}\lambda,\bar{\mu}\mu}(\vec{p}_{1},\vec{p}_{3},P)$ is defined as
@@
where
@@
By utilizing this effective interaction kernel and expanding the equation in terms of momenta and the bound energy, the effective Schrodinger-like equation Eq. (1) simplifies to the Schrodinger equation with the Coulomb potential. The contributions from interaction kernels involving higher orders of $\alpha_{e}$, $\vec{p}_{i}/m_{e,p}$, and $m_{e}/m_{p}$ beyond $\widetilde{K}^{(0)}(\vec{p}_{1},\vec{p}_{3},P)$ can be estimated systematically using the bound-state perturbative theory, order by order. This approach can also be applied in NRQED, where the power counting and analytical calculations are more straightforward.
@@ II.4 The form of the wave function
which results in the following equation:
@@ II.5 Numerical method
with $a=\frac{1}{\mu e^{2}}$ the Bohr radius.
@@
where $\Phi^{(0)}_{\alpha,\beta;i}(\vec{p}_{1})$ is simply $\Phi_{\alpha,\beta}(\vec{p}_{1})$ in Eq. (25) after replacing $\Phi_{nlm}(\vec{p}_{1})$ with $\phi^{(0)}_{il}(|\vec{p}_{1}|)Y_{lm}$. The integration over the angles $\Omega_{\vec{p}1}$ and $\Omega_{\vec{p}_{3}}$ can be performed analytically, while the integration over $|\vec{p_{1}}|$ and $|\vec{p}_{3}|$ can be done numerically with high precision. After diagonalizing this matrix, the energy spectrum $E_{n}\equiv P{0,n}-m_{e}-m_{p}$ is obtained approximately.
@@ III Numerical Results and discussion
In our numerical calculation, we set $n_{max}=100$, and the physical constants are taken as $m_{e}=0.510998950$ MeV, $m_{p}=938.27208816$ MeV, and $\alpha_{e}=\frac{1}{137.035999084}$. The relative precision of the numerical calculation for each matrix element reaches $10^{-20}$, ensuring that the absolute precision of the matrix elements is better than $10^{-18}$ eV. This guarantees the reliability of the numerical results. The numerical precision is also tested for the matrix elements involving the Coulomb and Breit potentials.
@@
$n$ peV
$\Delta E_{n,0}(S)$
$\Delta E_{n,0}(P)$
$\Delta E_{n,0}(D)$
$n=1$
-2349361
$n=2$
-289552
813
$n=3$
-85154
424
13
$n=4$
-35738
242
16
$n=5$
-18225
151
14
@@
$n$ peV
$\Delta E_{n,1}(S)$
$\Delta E_{n,1}(P)$
$\Delta E_{n,1}(D)$
$n=1$
-2349361
$n=2$
-289552
-44
$n=3$
-85154
50
3
$n=4$
-35738
48
6
$n=5$
-18225
37
6
@@
$n$ peV
$\Delta E_{n,2}(S)$
$\Delta E_{n,2}(P)$
$\Delta E_{n,2}(D)$
$n=1$
-2357360
$n=2$
-290555
-45
$n=3$
-85452
50
3
$n=4$
-35864
48
6
$n=5$
-18289
37
6
@@
$n$ peV
$\Delta E_{n}^{fin}(S)$
$\Delta E_{n}^{fin}(P)$
$\Delta E_{n}^{fin}(D)$
$n=1$
0
$n=2$
0
-857
$n=3$
0
-374
-10
$n=4$
0
-194
-10
$n=5$
0
-114
-8
@@
$n$ peV
$\Delta E_{n}^{hyf}(S)$
$\Delta E_{n}^{hyf}(P)$
$\Delta E_{n}^{hyf}(D)$
$n=1$
-7999
$n=2$
-1003
$-0.66$
$n=3$
-298
$-0.27$
$-0.007$
$n=4$
-126
$-0.14$
$-0.007$
$n=5$
-65
$-0.08$
$-0.005$
@@
In our calculation, we do not expand the OPE interaction kernel order by order in momenta, but instead solve the effective Schrodinger-like equation using numerical method. This approach enables us to include all contributions from ladder diagrams with the approximated propagator $\bar{G}_{0}$ and the full photon propagator, while excluding the crossed diagrams. Consequently, the comparison of our results with those obtained from bound-state perturbation theory is not straightforward.

### v3

@@ Abstract
In this work, we present expressions for the full effective potential corresponding to the one-photon exchange interaction within the framework of an effective Schrödinger-like equation, which is derived exactly from the Bethe-Salpeter equation in quantum electrodynamics. The final effective potential is expressed in terms of eight scalar functions. When these eight scalar functions are expanded in terms of velocities order by order, we can return to the non-relativistic effective potential systematically organized by velocities. By retaining the exact momentum dependence in the effective potential, we estimate its corrections to the energy spectrum of hydrogen using a highly precise numerical method. The comparison of the numerical results with those obtained using conventional bound-state perturbative theory is discussed. Our calculations suggest that it is possible to accurately account for all relativistic contributions using this method. It would be interesting to extend these calculations to positronium, muonic hydrogen, and scenarios involving nuclear structure and radiative corrections.
@@ I Introduction
In the bound-state perturbative theory, the interaction kernel in the effective Schrödinger-like equation or the effective Dirac-like equations is expanded order by order in terms of the fine structure constant $\alpha_{e}$, velocities $\vec{p}_{i}/m_{e,p}$, and $m_{e}/m_{p}$, where $\vec{p}_{i}$ represents the three momenta of the particles in the center-of-mass frame, and $m_{e}$ and $m_{p}$ denote the masses of the electron and proton, respectively. The expansion of the interaction kernel in terms of velocities implies that only contributions from low momenta are accurately considered, while contributions from high momenta need to be treated separately.
On the other hand, the wave functions derived from the Schrödinger-like equation with the Coulomb potential decrease polynomily on the momentum in the momentum space. These wave functions are used for the entire momentum range in the bound-state perturbative theory. The integration of the wave functions and interaction kernels encompasses contributions from both low and high momenta. This can introduce additional ultraviolet (UV) divergences, which differ somewhat from the usual UV divergences in the scattering amplitudes arising from loop integrals in radiative corrections. To handle these UV divergences, regularization methods are employed Pachucki-PRA-1997 ; Czarnecki-PRA-1999 ; Adkins-PRA-2020 . After considering all the contributions at the same order, these UV divergences eventually cancel each other out, and the final results are independent of the regularization procedure and are free of divergences.
Throughout history, various other methods have also been employed to investigate corrections beyond the Coulomb potential. These include the use of the Dirac equation with an effective potential Grotch-RMP-1969 , the external field approximation Eides-Book-2007 , the quasi-potential approach Faustov-1985-TMP ; Faustov-1999-zETF ; Martynenko-2005-JETP , and the Foldy-Wouthuysen transformation method Foldy-PR-1950 ; Pachucki-2005-PRA ; Zhou-JPB-2023 , among others.
In this work, we focus our discussion within the framework of the effective Schrödinger-like equation, which is derived exactly from the Bethe-Salpeter equation in QED Lepage-1978-PRA-Schrodinger-like-equation . Different from the usual bound-state perturbative theory applied in hydrogen-like system, we do not expand the interaction kernel in terms of velocities and $m_{e}/m_{p}$, but retain them in its relativistic form. The relativistic form effective potential has correct behavior at high energy or short distance. By taking the one-photon-exchange (OPE) interaction kernel as an example, we express the corresponding relativistic form effective potential by eight scalar functions which includes all relativistic corrections at the order of $\alpha_{e}$. This effective potential are valid across the entire momentum range and can reproduce the usual low energy behavior. By this potential, the additional UV divergences in the bound-state perturbative theory naturally disappear. Furthermore, we use a high precise numerical method calculate the energy spectrum by this potential. This numerical calculation can be regarded as a full result incorporating the OPE interaction, and it provides an interesting basis for comparison with the results obtained using the usual bound-state perturbative theory.
@@ II Basic Formula
II.1 The effective Schrödinger-like equation
In QED, the bound states of the $ep$ system in the center frame can be accurately described by the following effective Schrödinger-like equation Lepage-1978-PRA-Schrodinger-like-equation :
@@ II.1 The effective Schrödinger-like equation
where $P=(P_{0},\vec{0})\equiv p_{1}+p_{2}=p_{3}+p_{4}$, $p_{1,2,3,4}$ are the momenta of final and intermediate fermions as shown in Fig.1, $E_{i}^{(e)}\equiv\sqrt{\vec{p}_{i}^{2}+m_{e}^{2}}$, $E_{i}^{(p)}\equiv\sqrt{\vec{p}_{i}^{2}+m_{p}^{2}}$, and $\lambda,\mu$ are the indexes related with projected spinors. The interaction kernel $\widetilde{K}_{\bar{\lambda}\lambda,\bar{\mu}\mu}(\vec{p}_{1},\vec{p}_{3},P)$ is defined as
@@
where $\lambda,\bar{\lambda},\mu,\bar{mu}$ are the helicity of the projected spinors and
@@
By utilizing this effective interaction kernel and expanding the equation in terms of momenta and the bound energy, the effective Schrödinger-like equation Eq. (1) simplifies to the Schrödinger equation with the Coulomb potential. The contributions from interaction kernels involving higher orders of $\alpha_{e}$, $\vec{p}_{i}/m_{e,p}$, and $m_{e}/m_{p}$ beyond $\widetilde{K}^{(0)}(\vec{p}_{1},\vec{p}_{3},P)$ can be estimated systematically using the bound-state perturbative theory, order by order. This approach can also be applied in NRQED, where the power counting and analytical calculations are more straightforward.
@@ II.4 The form of the wave function
where $\chi$ is the Pauli spinor and $\bar{\lambda},\bar{\mu}$ are the spin of the Pauli spinors. This results in the following equation:
@@ II.5 Numerical method
with $C$ the Gegenbauer polynomial and $a=\frac{1}{\mu e^{2}}$ the Bohr radius.
@@
where $\Phi^{(0)}_{\alpha,\beta;i}(\vec{p}_{1})$ is simply $\Phi_{\alpha,\beta}(\vec{p}_{1})$ in Eq. (25) after replacing $\Phi_{nlm}(\vec{p}_{1})$ with $\phi^{(0)}_{il}(|\vec{p}_{1}|)Y_{lm}$. The integration over the angles $\Omega_{\vec{p}1}$ and $\Omega_{\vec{p}_{3}}$ can be performed analytically, while the integration over $|\vec{p_{1}}|$ and $|\vec{p}_{3}|$ can be done numerically with high precision. After diagonalizing this matrix, the energy spectrum $E_{n}\equiv P_{0,n}-m_{e}-m_{p}$ is obtained approximately.
We would like to highlight a significant difference between the method described above and the bound state perturbative theory. In the latter, non-relativistic potentials are utilized order by order. These non-relativistic potentials can lead to additional UV divergences in certain matrix elements, which are only cancelled when all intermediate states are considered. In contrast, our calculation correctly accounts for the higher energy behavior of the effective potential, and there are no additional UV divergences; a finite $n_{max}$ can yield precise results. We will elaborate on this property in detail in the next section.
@@ III Numerical Results and discussion
In our numerical calculation, we set $n_{max}=100$, and the physical constants are taken as $m_{e}=0.510998950$ MeV, $m_{p}=938.27208816$ MeV, and $137.0359990841/\alpha_{e}=$ 137.035999084. The relative precision of the numerical calculation for each matrix element reaches $10^{-20}$, ensuring that the absolute precision of the matrix elements is better than $10^{-18}$ eV. This guarantees the reliability of the numerical results. The numerical precision is also tested for the matrix elements involving the Coulomb and Breit potentials.
@@
$n$ peV $\Delta E_{n,0}(S)$ $\Delta E_{n,0}(P)$ $\Delta E_{n,0}(D)$
$n=1$2349361 - -
$n=2$289552 813 -
$n=3$85154 424 13
$n=4$35738 242 16
$n=5$18225 151 14
@@
$n$ peV $\Delta E_{n,1}(S)$ $\Delta E_{n,1}(P)$ $\Delta E_{n,1}(D)$
$n=1$2349361 - -
$n=2$289552 -44 -
$n=3$85154 50 3
$n=4$35738 48 6
$n=5$18225 37 6
@@
$n$ peV $\Delta E_{n,2}(S)$ $\Delta E_{n,2}(P)$ $\Delta E_{n,2}(D)$
$n=1$2357360 - -
$n=2$290555 -45 -
$n=3$85452 50 3
$n=4$35864 48 6
$n=5$18289 37 6
@@
$n$ peV $\Delta E_{n}^{fin}(S)$ $\Delta E_{n}^{fin}(P)$ $\Delta E_{n}^{fin}(D)$
$n=1$ 0 - -
$n=2$ 0 -857 -
$n=3$ 0 -374 -10
$n=4$ 0 -194 -10
$n=5$ 0 -114 -8
@@
$n$ peV $\Delta E_{n}^{hyf}(S)$ $\Delta E_{n}^{hyf}(P)$ $\Delta E_{n}^{hyf}(D)$
$n=1$7999 - -
$n=2$1003 $-0.66$
$n=3$298 $-0.27$ $-0.007$
$n=4$126 $-0.14$ $-0.007$
$n=5$65 $-0.08$ $-0.005$
@@
In our calculation, we do not expand the OPE interaction kernel order by order in momenta, but instead solve the effective Schrödinger-like equation using numerical method. This approach enables us to include all contributions from ladder diagrams with the approximated propagator $\bar{G}_{0}$ and the full photon propagator, while excluding the crossed diagrams. Consequently, the comparison of our results with those obtained from bound-state perturbation theory is not straightforward.